if triangle abc is similar to traingle zxy, the measure of angle a is 50 and the measure of angle c is 30, what is the measure of angle x?
Answer by jim_thompson5910(28598)

The measure of angle b is 180-(a+c) = 180-(50+30) = 180-80 = 100 (remember that all 3 angles in a triangle add to 180 degrees)


Now recall that similar triangles have equal corresponding angles. Since angle b corresponds to angle x (draw a picture if you can't see this), this means that the measure of angle x is 100 degrees.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim